OpenStudy (anonymous):

Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? Some help please!!!

OpenStudy (anonymous):

\[f(x)=\frac{3x-4}{5}\]?

OpenStudy (anonymous):

Yes! Please help me i am so confused:c

OpenStudy (campbell_st):

well there are no answer choices listed to match to question

OpenStudy (anonymous):

My bad the choices are: A.) f−1(x) = The quantity of 5x plus 4, divided by 3. B.) f−1(x) = The quantity of 5x minus 4, divided by 3. C.) f−1(x) = The quantity of negative 3x minus 4, divided by 5. D.) f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (campbell_st):

so you are being asked to find the inverse of the function...?

OpenStudy (anonymous):

Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? that is all it says...

OpenStudy (campbell_st):

so do the answer choices start with \[f^{-1} (x)\]

OpenStudy (campbell_st):

ok... I think they question is based around you being able to find the inverse function. so to do that.. you're original function can be written as \[y = \frac{3x -4}{5}\] to find the inverse function, swap x and y \[x = \frac{3y - 4}{5}\] the task is to may y the subject.... can you do that..?

OpenStudy (campbell_st):

so multiply both sides by 5 5x = 3y - 4 add 4 to both sides 5x + 4 = 3y divide both sides by 3 \[y = \frac{5x + 4}{3}\] so the inverse function is written as \[f^{-1}(x) = \frac{5x + 4}{3}\]

OpenStudy (anonymous):

Honestly you're amazing thank you so much i actually get it! Do you think you could help me with some others i'm struggling with?